# Linear Discriminant Regularized Regression

### v2

@@ B.4 Proof of 4
${t_{\ell}\over C}=\sqrt{\Delta_{\infty}}\sqrt{L\log(n)\over n}+\|\Sigma_{w}^{1/2}(\widehat{B}^{*}_{\ell}-B^{*}_{\ell})\|_{2}\sqrt{\Delta_{\infty}+\log(n)}+|(\widehat{\mu}_{\ell}-\mu_{\ell})^{\top}(\widehat{B}^{*}_{\ell}-B^{*}_{\ell})|.$
@@
$\mathbb{E}_{{\bm{D}}}[\mathcal{R}(\widehat{g})]$ $\leq\mathbb{E}_{{\bm{D}}}\left[\min\left\{C\sum_{\ell=1}^{L}t_{\ell}^{2}+L\sum_{\ell=1}^{L}\mathbb{P}\left\{|\widehat{G}_{\ell}(X)-G_{\ell}(X)|\geq t_{\ell}\mid{\bm{D}}\right\},~{}1\right\}1\{\mathcal{E}_{\mu}\cap\mathcal{E}_{\pi}\}\right]$
@@
$\lesssim\mathbb{E}_{{\bm{D}}}\left[\min\left\{\Delta_{\infty}{L^{2}\log(n)\over n}+\mathcal{Q}(\widehat{B}^{*})+{L\over n},~{}1\right\}1\{\mathcal{E}_{\mu}\cap\mathcal{E}_{\pi}\}\right]+{1\over n}$
$\lesssim\mathbb{E}_{{\bm{D}}}\left[\min\left\{\Delta_{\infty}{L^{2}\log(n)\over n}+\mathcal{Q}(\widehat{B}^{*}),~{}1\right\}\right],$
@@ B.5 Proof of Theorem 5
$\delta=C\sqrt{\log(n)\over nL}+C{\omega_{2}\over\sqrt{L}}$ (B.10)
@@
$\lambda_{K}(\widehat{H})$ $\geq\lambda_{K}(H)-\|\widehat{H}-H\|_{\rm op}$ by Weyl’s inequality
$=\left[\lambda_{1}(\Omega)\right]^{-1}-\|\widehat{H}-H\|_{\rm op}$ by $\text{by }\Omega=H^{-1}$
@@
$\|\Sigma_{w}^{1/2}(\widehat{B}^{*}-B^{*})\|_{F}^{2}~{}\lesssim~{}\left(\omega_{2}^{2}+\delta^{2}L\Delta_{\infty}\right)L^{2}\Delta_{\infty}^{2},$ (B.11)
$\sum_{\ell=1}^{L}\left[(\widehat{\mu}_{\ell}-\mu_{\ell})^{\top}(\widehat{B}^{*}_{\ell}-B^{*}_{\ell})\right]^{2}\lesssim\left(\omega_{1}^{2}{\log(n\vee p)\over n}+\delta^{2}L\Delta_{\infty}{\log(n)\over n}\right)L^{3}\Delta_{\infty}^{2}$ (B.12)
which, in conjunction with (B.10), $\Delta\asymp 1$, $L\log(n)\leq cn$ and $\omega_{2}\sqrt{L}\leq c$, yield the claim.
To prove (B.11), we write $\widehat{\Omega}=\widehat{H}^{-1}$ for simplicity. By definition,
@@
$=(\widehat{B}-B)\widehat{\Omega}+B^{*}(H\widehat{\Omega}-\mathbf{I}_{L})$ by $\text{by }H\Omega=\mathbf{I}_{L}$
$=(\widehat{B}-B)\widehat{\Omega}+B^{*}(H-\widehat{H})\widehat{\Omega}$ by $\text{by }\widehat{H}\widehat{\Omega}=\mathbf{I}_{L}.$ (B.13)
@@
$\|\widehat{\Omega}\|_{\rm op}\leq 2\|\Omega\|_{\rm op}\lesssim L\Delta_{\infty}.$ (B.14)
It now follows from 2 that
@@
$\leq\left(\omega_{2}+\delta\sqrt{L\Delta_{\infty}}\right)L\Delta_{\infty},$ (B.15)
proving (B.11).
We now prove (B.12). Using the identity (B.5), we have, on the event $\mathcal{E}_{H}$,
@@
$\leq 2\sum_{\ell=1}^{L}\left[(\widehat{\mu}_{\ell}-\mu_{\ell})^{\top}(\widehat{B}-B)\widehat{\Omega}_{\ell}\right]^{2}+2\sum_{\ell=1}^{L}\left[(\widehat{\mu}_{\ell}-\mu_{\ell})^{\top}B^{*}(H-\widehat{H})\widehat{\Omega}_{\ell}\right]^{2}.$ (B.16)
@@
$\max_{\ell\in[L]}\|\widehat{\mu}_{\ell}-\mu_{\ell}\|_{\infty}^{2}\lesssim{\|\Sigma_{w}\|_{\infty}}{L\log(n\vee p)\over n}$ (B.17)
@@
By invoking 2 and applying inequalities (B.14) and (B.17), we obtain
$\sum_{\ell=1}^{L}\left[(\widehat{\mu}_{\ell}-\mu_{\ell})^{\top}(\widehat{B}^{*}_{\ell}-B^{*}_{\ell})\right]^{2}\lesssim\omega_{1}^{2}{\|\Sigma_{w}\|_{\infty}}{\log(n\vee p)\over n}L^{3}\Delta_{\infty}^{2}$ (B.18)
with probability at least $1-n^{-2}$. Regarding the second term on the right in (B.5), we have
@@
on the event $\mathcal{E}_{H}$. Further invoking (B.7) and (B.14) yields that
$\sum_{\ell=1}^{L}\left[(\widehat{\mu}_{\ell}-\mu_{\ell})^{\top}B^{*}(H-\widehat{H})\widehat{\Omega}_{\ell}\right]^{2}$ $~{}\lesssim\delta^{2}{\log(n)\over n}L^{4}\Delta_{\infty}^{3}$ (B.19)
with probability at least $1-n^{-2}$. Combining (B.18) with (B.19) proves (B.12), and hence completes the proof. ∎
@@
$\begin{split}\widehat{H}-H&=D_{\widehat{\pi}}-D_{\pi}-\left(\widehat{B}^{\top}\widehat{\Sigma}\widehat{B}-B^{\top}\Sigma B\right)\\ &=D_{\widehat{\pi}}-D_{\pi}-B^{\top}(\widehat{\Sigma}-\Sigma)B-(\widehat{B}-B)^{\top}\widehat{\Sigma}(\widehat{B}-B)\\ &\qquad-(\widehat{B}-B)^{\top}\widehat{\Sigma}B-B^{\top}\widehat{\Sigma}(\widehat{B}-B).\end{split}$ (B.20)
@@
$\|D_{\widehat{\pi}}-D_{\pi}\|_{\rm op}\leq\max_{\ell\in[L]}|\widehat{\pi}_{\ell}-\pi_{\ell}|\lesssim\sqrt{\log(n)\over nL}.$ (B.21)
Regarding the second term, 2 ensures that
$\|(\widehat{B}-B)^{\top}\widehat{\Sigma}(\widehat{B}-B)\|_{\rm op}\leq\|\widehat{\Sigma}^{1/2}(\widehat{B}-B)\|_{F}^{2}\leq\omega_{2}^{2}.$ (B.22)
@@
$\|B^{\top}(\widehat{\Sigma}-\Sigma)B\|_{\rm op}\lesssim{\Delta_{\rm op}\over L+\Delta_{\rm op}}\sqrt{\Delta_{\infty}\log(n)\over nL}+{\Delta_{\rm op}\over L+\Delta_{\rm op}}{\Delta_{\infty}\log(n)\over n}.$ (B.23)
@@
deduced from Lemma 14. Combining the above bounds with (B.21), (B.22) and (B.23) gives
@@
B.6 Proof of Theorem 6: the lasso estimator of $B$
@@
$\begin{split}&\max_{\ell\in[L]}{1\over\sqrt{n}}\|{\mathbf{X}}(\widehat{B}_{\ell}-B_{\ell})\|_{2}~{}\lesssim~{}\kappa_{s}^{-1/2}\lambda\sqrt{s}\\ &\max_{\ell\in[L]}\|\Sigma^{1/2}(\widehat{B}_{\ell}-B_{\ell})\|_{2}~{}\lesssim~{}\kappa_{s}^{-1/2}\lambda\sqrt{s}\\ &\max_{\ell\in[L]}\|\widehat{B}_{\ell}-B_{\ell}\|_{2}~{}\lesssim~{}\kappa_{s}^{-1}\lambda\sqrt{s}\end{split}$ (B.24)
@@
$\mathcal{E}_{\lambda}:=\left\{{1\over n}\|{\mathbf{X}}^{\top}({\mathbf{Y}}-{\mathbf{X}}B)\|_{\infty}\leq{1\over 2}\lambda\right\}$ (B.25)
@@
$\mathcal{E}_{RE}:=\left\{\min_{S\subseteq[p],|S|\leq s}\min_{v\in\mathcal{C}(S,3)}{\|{\mathbf{X}}v\|_{2}^{2}\over n\|v\|_{2}^{2}}\geq{\kappa_{s}\over 2}\right\}.$ (B.26)
@@
$\Gamma_{s}:=\bigcup_{S\subseteq[p]:\ |S|=\lfloor Cs\rfloor}E_{S}$ (B.27)
@@
${9\over 10}\|\Sigma^{1/2}v\|_{2}\leq\|\widehat{\Sigma}^{1/2}v\|_{2}\leq{11\over 10}\|\Sigma^{1/2}v\|_{2}\qquad\forall~{}v\in\Gamma_{s}$ (B.28)
holds with probability at least $1-(n\vee p)^{-2}$. According to Rudelson and Zhou (2012, Theorem 3), (B.28) implies that there exists $S\subseteq[p]$ with $|S|\leq s$ such that
$u^{\top}\widehat{\Sigma}u\geq{1\over 2}u^{\top}\Sigma u$ (B.29)
holds for any $u\in\mathcal{C}(S,3)$. Eq. B.29 in conjunction with 5 ensures that $\mathcal{E}_{RE}$ holds.
It is now relatively straightforward to finish the proof of (B.24). Standard arguments (Bühlmann and Van de Geer, 2011; Giraud, 2021) yield that on the event $\mathcal{E}_{\lambda}$, for any $\ell\in[L]$,
@@
$\widehat{B}_{\ell}-B_{\ell}\in\mathcal{C}(S_{\ell},3)$ (B.30)
with $S_{\ell}:={\rm supp}(B_{\ell})$. Invoking (B.29), we find on the event $\mathcal{E}_{RE}$
@@
B.6.1 Two key lemmas used in the proof of Theorem 6
@@
${1\over n}{\mathbf{X}}^{\top}({\mathbf{Y}}-{\mathbf{X}}B)$ $={1\over n}{\mathbf{X}}^{\top}{\mathbf{Y}}-\Sigma_{XY}-\widehat{\Sigma}B+\Sigma B$ (B.31)
We first analyze the term $(\widehat{\Sigma}-\Sigma)B$. Pick any $j\in[p]$, $\ell\in[L]$ and fix $t\geq 0$. We observe from Lemma 21 and the inequality $L\leq n$ that $\mathcal{E}_{\pi}$ holds with probability at least $1-2n^{-1}$. Next, we invoke Lemma 18 with $v_{1}={\bm{e}}_{j}$, $v_{2}=B_{\ell}$, $t=C\log(n\vee p)$ and we use (C.5), (C.6) and (C.7) in Lemma 14 to derive that the inequalities
@@
$\max_{j\in[p],\ \ell\in[L]}\left|{\bm{e}}_{j}^{\top}(\widehat{\Sigma}-\Sigma)B_{\ell}\right|$ $\lesssim\sqrt{\|\Sigma_{w}\|_{\infty}+\|{M}\|_{\infty}^{2}}\sqrt{\log(n\vee p)\over nL}$ (B.32)
@@
To bound the first term in (B.31), by recalling that $\Sigma_{XY}={M}D_{\pi}$, we find
@@
$\max_{j\in[p],\ \ell\in[L]}\left|{1\over n}{\mathbf{X}}_{j}^{\top}{\mathbf{Y}}_{\ell}-{\bm{e}}_{j}^{\top}\Sigma_{XY}{\bm{e}}_{\ell}\right|\lesssim\sqrt{\|\Sigma_{w}\|_{\infty}\log(p\vee n)\over nL}+\|{M}\|_{\infty}\sqrt{\log(n)\over nL}$ (B.33)
with probability $1-n^{-1}$. Combining (B.32) and (B.33) completes the proof. ∎
@@
$\sup_{v\in\Gamma_{s}^{\prime}}v^{\top}(\widehat{\Sigma}-\Sigma)v\leq 3\max_{u\in\mathcal{N}}u^{\top}(\widehat{\Sigma}-\Sigma)u.$ (B.34)
@@
We will bound the right hand side (B.34) via the same arguments used to prove Lemma 18, except for the control of the term $\sum_{k}(\widehat{\pi}_{k}-\pi_{k})v^{\top}\mu_{k}\mu_{k}^{\top}v$. Specifically, repeating the proof of Lemma 18 with $v_{1}=v_{2}=v$, for any $v\in\mathcal{N}$, gives that for any $t\geq 0$
@@
holds uniformly over $\mathcal{N}$. The proof is complete in view of (B.34). ∎
B.7 Proof of Theorem 8: the reduced-rank estimator of $B$

### v3

@@ A.3 Proof of 4
${t_{\ell}\over C}=\sqrt{\Delta}\sqrt{L\log(n)\over n}+\|\Sigma_{w}^{1/2}(\widehat{B}^{*}_{\ell}-B^{*}_{\ell})\|_{2}\sqrt{\Delta+\log(n)}+|(\widehat{\mu}_{\ell}-\mu_{\ell})^{\top}(\widehat{B}^{*}_{\ell}-B^{*}_{\ell})|.$
@@
[…]
@@
$\lesssim\mathbb{E}_{{\bm{D}}}\left[\min\left\{\Delta{L^{2}\log(n)\over n}+\mathcal{Q}(\widehat{B}^{*})+{L\over n},~1\right\}1\{\mathcal{E}_{\mu}\cap\mathcal{E}_{\pi}\}\right]+{1\over n}$
$\lesssim\mathbb{E}_{{\bm{D}}}\left[\min\left\{\Delta{L^{2}\log(n)\over n}+\mathcal{Q}(\widehat{B}^{*}),~1\right\}\right],$
@@ A.4 Proof of Theorem 5
$\delta=C\sqrt{\log(n)\over nL}+C{\omega_{2}\over\sqrt{L}}$ (A.10)
@@
[…]
$=\left[\lambda_{1}(\Omega)\right]^{-1}-\|\widehat{H}-H\|_{\rm op}$ $\text{by }\Omega=H^{-1}$
@@
$\|\Sigma_{w}^{1/2}(\widehat{B}^{*}-B^{*})\|_{F}^{2}~\lesssim~\left(\omega_{2}^{2}+\delta^{2}L\Delta_{\infty}\right)L^{2}\Delta_{\infty}^{2},$ (A.11)
$\sum_{\ell=1}^{L}\left[(\widehat{\mu}_{\ell}-\mu_{\ell})^{\top}(\widehat{B}^{*}_{\ell}-B^{*}_{\ell})\right]^{2}\lesssim\left(\omega_{1}^{2}{\log(n\vee p)\over n}+\delta^{2}L\Delta_{\infty}{\log(n)\over n}\right)L^{3}\Delta_{\infty}^{2}$ (A.12)
which, in conjunction with (A.10), $\Delta\asymp 1$, $L\log(n)\leq cn$ and $\omega_{2}\sqrt{L}\leq c$, yield the claim.
To prove (A.11), we write $\widehat{\Omega}=\widehat{H}^{-1}$ for simplicity. By definition,
@@
$=(\widehat{B}-B)\widehat{\Omega}+B^{*}(H\widehat{\Omega}-\mathbf{I}_{L})$ $\text{by }H\Omega=\mathbf{I}_{L}$
$=(\widehat{B}-B)\widehat{\Omega}+B^{*}(H-\widehat{H})\widehat{\Omega}$ $\text{by }\widehat{H}\widehat{\Omega}=\mathbf{I}_{L}.$ (A.13)
@@
$\|\widehat{\Omega}\|_{\rm op}\leq 2\|\Omega\|_{\rm op}\lesssim L\Delta_{\infty}.$ (A.14)
It now follows from 1 that
@@
$\leq\left(\omega_{2}+\delta\sqrt{L\Delta_{\infty}}\right)L\Delta_{\infty},$ (A.15)
proving (A.11).
We now prove (A.12). Using the identity (A.4), we have, on the event $\mathcal{E}_{H}$,
@@
$\leq 2\sum_{\ell=1}^{L}\left[(\widehat{\mu}_{\ell}-\mu_{\ell})^{\top}(\widehat{B}-B)\widehat{\Omega}_{\ell}\right]^{2}+2\sum_{\ell=1}^{L}\left[(\widehat{\mu}_{\ell}-\mu_{\ell})^{\top}B^{*}(H-\widehat{H})\widehat{\Omega}_{\ell}\right]^{2}.$ (A.16)
@@
$\max_{\ell\in[L]}\|\widehat{\mu}_{\ell}-\mu_{\ell}\|_{\infty}^{2}\lesssim{\|\Sigma_{w}\|_{\infty}}{L\log(n\vee p)\over n}$ (A.17)
@@
By invoking 1 and applying inequalities (A.14) and (A.17), we obtain
$\sum_{\ell=1}^{L}\left[(\widehat{\mu}_{\ell}-\mu_{\ell})^{\top}(\widehat{B}^{*}_{\ell}-B^{*}_{\ell})\right]^{2}\lesssim\omega_{1}^{2}{\|\Sigma_{w}\|_{\infty}}{\log(n\vee p)\over n}L^{3}\Delta_{\infty}^{2}$ (A.18)
with probability at least $1-n^{-2}$. Regarding the second term on the right in (A.4), we have
@@
on the event $\mathcal{E}_{H}$. Further invoking (A.7) and (A.14) yields that
$\sum_{\ell=1}^{L}\left[(\widehat{\mu}_{\ell}-\mu_{\ell})^{\top}B^{*}(H-\widehat{H})\widehat{\Omega}_{\ell}\right]^{2}$ $~\lesssim\delta^{2}{\log(n)\over n}L^{4}\Delta_{\infty}^{3}$ (A.19)
with probability at least $1-n^{-2}$. Recall that $\Delta_{\infty}\leq C\Delta$ from 2. Combining (A.18) with (A.19) proves (A.12), and hence completes the proof. ∎
@@
$\begin{split}\widehat{H}-H&=D_{\widehat{\pi}}-D_{\pi}-\left(\widehat{B}^{\top}\widehat{\Sigma}\widehat{B}-B^{\top}\Sigma B\right)\\ &=D_{\widehat{\pi}}-D_{\pi}-B^{\top}(\widehat{\Sigma}-\Sigma)B-(\widehat{B}-B)^{\top}\widehat{\Sigma}(\widehat{B}-B)\\ &\qquad-(\widehat{B}-B)^{\top}\widehat{\Sigma}B-B^{\top}\widehat{\Sigma}(\widehat{B}-B).\end{split}$ (A.20)
@@
$\|D_{\widehat{\pi}}-D_{\pi}\|_{\rm op}\leq\max_{\ell\in[L]}|\widehat{\pi}_{\ell}-\pi_{\ell}|\lesssim\sqrt{\log(n)\over nL}.$ (A.21)
Regarding the second term, 1 ensures that
$\|(\widehat{B}-B)^{\top}\widehat{\Sigma}(\widehat{B}-B)\|_{\rm op}\leq\|\widehat{\Sigma}^{1/2}(\widehat{B}-B)\|_{F}^{2}\leq\omega_{2}^{2}.$ (A.22)
@@
$\|B^{\top}(\widehat{\Sigma}-\Sigma)B\|_{\rm op}\lesssim{\Delta_{\rm op}\over L+\Delta_{\rm op}}\sqrt{\Delta_{\infty}\log(n)\over nL}+{\Delta_{\rm op}\over L+\Delta_{\rm op}}{\Delta_{\infty}\log(n)\over n}.$ (A.23)
@@
deduced from Lemma 14. Combining the above bounds with (A.21), (A.22) and (A.23) gives
@@
A.5 Proof of Theorem 6: the lasso estimator of $B$
@@
$\begin{split}&\max_{\ell\in[L]}{1\over\sqrt{n}}\|{\mathbf{X}}(\widehat{B}_{\ell}-B_{\ell})\|_{2}~\lesssim~\kappa_{s}^{-1/2}\lambda\sqrt{s}\\ &\max_{\ell\in[L]}\|\Sigma^{1/2}(\widehat{B}_{\ell}-B_{\ell})\|_{2}~\lesssim~\kappa_{s}^{-1/2}\lambda\sqrt{s}\\ &\max_{\ell\in[L]}\|\widehat{B}_{\ell}-B_{\ell}\|_{2}~\lesssim~\kappa_{s}^{-1}\lambda\sqrt{s}\end{split}$ (A.24)
@@
$\mathcal{E}_{\lambda}:=\left\{{1\over n}\|{\mathbf{X}}^{\top}({\mathbf{Y}}-{\mathbf{X}}B)\|_{\infty}\leq{1\over 2}\lambda\right\}$ (A.25)
@@
$\mathcal{E}_{RE}:=\left\{\min_{S\subseteq[p],|S|\leq s}\min_{v\in\mathcal{C}(S,3)}{\|{\mathbf{X}}v\|_{2}^{2}\over n\|v\|_{2}^{2}}\geq{\kappa_{s}\over 2}\right\}.$ (A.26)
@@
$\Gamma_{s}:=\bigcup_{S\subseteq[p]:\ |S|=\lfloor Cs\rfloor}E_{S}$ (A.27)
@@
${9\over 10}\|\Sigma^{1/2}v\|_{2}\leq\|\widehat{\Sigma}^{1/2}v\|_{2}\leq{11\over 10}\|\Sigma^{1/2}v\|_{2}\qquad\forall~v\in\Gamma_{s}$ (A.28)
holds with probability at least $1-(n\vee p)^{-2}$. According to Rudelson and Zhou (2012, Theorem 3), (A.28) implies that there exists $S\subseteq[p]$ with $|S|\leq s$ such that
$u^{\top}\widehat{\Sigma}u\geq{1\over 2}u^{\top}\Sigma u$ (A.29)
holds for any $u\in\mathcal{C}(S,3)$. Eq. A.29 in conjunction with 5 ensures that $\mathcal{E}_{RE}$ holds.
It is now relatively straightforward to finish the proof of (A.24). Standard arguments (Bühlmann and Van de Geer, 2011; Giraud, 2021) yield that on the event $\mathcal{E}_{\lambda}$, for any $\ell\in[L]$,
@@
$\widehat{B}_{\ell}-B_{\ell}\in\mathcal{C}(S_{\ell},3)$ (A.30)
with $S_{\ell}:={\rm supp}(B_{\ell})$. Invoking (A.29), we find on the event $\mathcal{E}_{RE}$
@@
A.5.1 Two key lemmas used in the proof of Theorem 6
@@
${1\over n}{\mathbf{X}}^{\top}({\mathbf{Y}}-{\mathbf{X}}B)$ $={1\over n}{\mathbf{X}}^{\top}{\mathbf{Y}}-\Sigma_{XY}-\widehat{\Sigma}B+\Sigma B$ (A.31)
We first analyze the term $(\widehat{\Sigma}-\Sigma)B$. Pick any $j\in[p]$, $\ell\in[L]$ and fix $t\geq 0$. We observe from Lemma 21 and the inequality $L\leq n$ that $\mathcal{E}_{\pi}$ holds with probability at least $1-2n^{-1}$. Next, we invoke Lemma 18 with $v_{1}={\bm{e}}_{j}$, $v_{2}=B_{\ell}$, $t=C\log(n\vee p)$ and we use (B.5), (B.6) and (B.7) in Lemma 14 to derive that the inequalities
@@
$\max_{j\in[p],\ \ell\in[L]}\left|{\bm{e}}_{j}^{\top}(\widehat{\Sigma}-\Sigma)B_{\ell}\right|$ $\lesssim\sqrt{\|\Sigma_{w}\|_{\infty}+\|{M}\|_{\infty}^{2}}\sqrt{\log(n\vee p)\over nL}$ (A.32)
@@
To bound the first term in (A.31), by recalling that $\Sigma_{XY}={M}D_{\pi}$, we find
@@
$\max_{j\in[p],\ \ell\in[L]}\left|{1\over n}{\mathbf{X}}_{j}^{\top}{\mathbf{Y}}_{\ell}-{\bm{e}}_{j}^{\top}\Sigma_{XY}{\bm{e}}_{\ell}\right|\lesssim\sqrt{\|\Sigma_{w}\|_{\infty}\log(p\vee n)\over nL}+\|{M}\|_{\infty}\sqrt{\log(n)\over nL}$ (A.33)
with probability $1-n^{-1}$. Combining (A.32) and (A.33) completes the proof. ∎
@@
$\sup_{v\in\Gamma_{s}^{\prime}}v^{\top}(\widehat{\Sigma}-\Sigma)v\leq 3\max_{u\in\mathcal{N}}u^{\top}(\widehat{\Sigma}-\Sigma)u.$ (A.34)
@@
We will bound the right hand side (A.34) via the same arguments used to prove Lemma 18, except for the control of the term $\sum_{k}(\widehat{\pi}_{k}-\pi_{k})v^{\top}\mu_{k}\mu_{k}^{\top}v$. Specifically, repeating the proof of Lemma 18 with $v_{1}=v_{2}=v$, for any $v\in\mathcal{N}$, gives that for any $t\geq 0$
@@
holds uniformly over $\mathcal{N}$. The proof is complete in view of (A.34). ∎
A.6 Proof of Theorem 8: the reduced-rank estimator of $B$